OpenStudy (natsent06):

if m= -64 and n= -6, find the value for each expression: _____ ∛-m and n⁵

OpenStudy (anonymous):

\[\sqrt[3]{-m}=\sqrt[3]{-(-64)}=\sqrt[3]{64}=..\] \[n=-6=>n^{5}=(-6)^{5}=-....\]

OpenStudy (natsent06):

?

OpenStudy (anonymous):

use the calculator for the final answers

OpenStudy (anonymous):

(-6)^5=(-6)*(-6)*(-6)*(-6)*(-6)

OpenStudy (natsent06):

4 and -7776

OpenStudy (natsent06):

?

OpenStudy (anonymous):

yup